# Tensor Product Attention Is All You Need

### v3

@@ E.1 Multi-Query Attention (MQA)
Multi-Query Attention (MQA) (Shazeer, 2019) significantly reduces memory usage by sharing keys and values across heads, while still preserving unique query projections. For a sequence of embeddings $\mathbf{X}\in\mathbb{R}^{T\times d_{\text{model}}}$,
@@
Hence, each head $i$ only has a distinct query $\mathbf{Q}_{i}\in\mathbb{R}^{T\times d_{h}}$, but shares the same key $\mathbf{K}_{\text{shared}}\in\mathbb{R}^{T\times d_{h}}$ and value $\mathbf{V}_{\text{shared}}\in\mathbb{R}^{T\times d_{h}}$. In practice, this means:
@@
By sharing these key and value projections, MQA cuts down on memory usage (especially for the key-value cache in autoregressive inference) but loses some expressivity since all heads must rely on the same key/value representations.
E.2 Grouped Query Attention (GQA)
Grouped Query Attention (GQA) (Ainslie et al., 2023) generalizes MHA and MQA by grouping heads. Specifically, we partition the $h$ total heads into $G$ groups. Each group has a single set of keys and values, but each individual head within that group still retains its own query projection. Formally, if $g(i)$ maps a head $i\in[h]$ to its group index $g\in[G]$, then:
@@ E.2 Grouped Query Attention (GQA)
Again, $\bm{W}^{K}_{g},\bm{W}^{V}_{g}\in\mathbb{R}^{d_{\text{model}}\times d_{h}}$ for each group $g$, and $\bm{W}^{Q}_{i}\in\mathbb{R}^{\,d_{\text{model}}\times d_{h}}$ for each head $i$. The complete output is again a concatenation of all heads:
@@
By adjusting $G$ between $1$ and $h$, GQA can interpolate between sharing all key/value projections across heads (i.e., MQA) and having one set of projections per head (i.e., MHA).
E.3 Multi-head Latent Attention (MLA)
Below, we briefly outline the Multi-head Latent Attention (MLA) approach used by DeepSeek-V2 (Liu et al., 2024a) and DeepSeek-V3 (Liu et al., 2024b). MLA introduces a low-rank compression of the keys and values to reduce the Key-Value (KV) caching cost at inference.
@@ E.3 Multi-head Latent Attention (MLA)
where $\bm{W}^{DKV}\in\mathbb{R}^{\,d_{\text{model}}\times d_{c}},\bm{W}^{UK}\in\mathbb{R}^{d_{c}\times d_{h}h},\bm{W}^{KR}\in\mathbb{R}^{d_{\text{model}}\times d_{h}^{R}},\bm{W}^{UV}\in\mathbb{R}^{d_{c}\times d_{h}h}$, and $\mathbf{C}^{KV}\in\mathbb{R}^{T\times d_{c}}$ is the compressed KV latent (with $d_{c}\ll d_{h}h$), and $\operatorname{RoPE}(\cdot)$ represents the RoPE transform applied to the separate key embeddings $\mathbf{K}^{R}$ of dimension $d_{h}^{R}$. Thus, only $\mathbf{C}^{KV}$ and $\mathbf{K}^{R}$ need to be cached, reducing KV memory usage while largely preserving performance compared to standard MHA (Vaswani et al., 2017).
@@
where $\bm{W}^{DQ}\in\mathbb{R}^{\,d_{\text{model}}\times d_{c}^{\prime}},\bm{W}^{UQ}\in\mathbb{R}^{\,d_{c}^{\prime}\times d_{h}h},\bm{W}^{QR}\in\mathbb{R}^{d_{c}^{\prime}\times d_{h}^{R}h}$. Here, $\mathbf{C}^{Q}\in\mathbb{R}^{T\times d_{c}^{\prime}}$ (with $d_{c}^{\prime}\ll d_{h}h$) is the compressed query latent. As above, each $\bm{W}^{DQ}$, $\bm{W}^{UQ}$, and $\bm{W}^{QR}$ connects these lower-dimensional query latents back to $h$ heads of dimension $d_{h}+d_{h}^{R}$.
@@
where $\bm{W}^{O}\in\mathbb{R}^{(d_{h}h)\times\,d_{\text{model}}}$ is the output projection.
In inference time, $\mathbf{C}^{KV}$ and $\mathbf{K}^{R}$ can be cached to accelerate decoding. In detail, when RoPE is ignored, the inner product $\mathbf{q}_{t,i}^{\top}\mathbf{k}_{s,i}$ (where $\mathbf{q}_{t,i},\mathbf{k}_{s,i}\in\mathbb{R}^{d}$) of the $i$-th head between $t$-th and $s$-th tokens can be calculated using the hidden state $\mathbf{x}_{t}\in\mathbb{R}^{d_{\text{model}}}$ for $t$-th token and the cached latent state $\mathbf{c}_{s}^{KV}\in\mathbb{R}^{d_{c}}$ for $s$-th token:
$\mathbf{q}_{t,i}^{\top}\mathbf{k}_{s,i}$ $=[(\bm{W}^{UQ}_{i})^{\top}(\bm{W}^{DQ}_{i})^{\top}\mathbf{x}_{t}]^{\top}[(\bm{W}^{UK}_{i})^{\top}\mathbf{c}^{KV}_{s}]$ (E.1)
$=\mathbf{x}_{t}^{\top}[\bm{W}^{DQ}_{i}\bm{W}^{UQ}_{i}(\bm{W}^{UK}_{i})^{\top}]\mathbf{c}^{KV}_{s},$ (E.2)
where $\bm{W}_{i}^{(\cdot)}$ is the $i$-th head of the original weight, and $[\bm{W}^{DQ}_{i}\bm{W}^{UQ}_{i}(\bm{W}^{UK}_{i})^{\top}]$ can be computed previously for faster decoding. However, this process fails when RoPE is considered according to Su (2024). Since RoPE can be considered as multiplication with a block-diagonal matrix $\mathbf{T}_{t}\in\mathbb{R}^{d_{h}\times d_{h}}$ (see Section 2.3), with the property (2.1) that $\mathbf{T}_{t}\mathbf{T}_{s}^{\top}=\mathbf{T}_{t-s}$, then
@@
Different from (E.2), acceleration by pre-computing $[\bm{W}^{DQ}_{i}\bm{W}^{UQ}_{i}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{T}_{t-s}}(\bm{W}^{UK}_{i})^{\top}]$ fails since it varies for different $(t,s)$ position pairs. Therefore, MLA adds the additional $\mathbf{k}_{t}^{R}$ part with a relatively smaller size for RoPE compatibility. In Section 3.2, we will show that TPA addresses the issue of RoPE-incompatibility by applying tensor product.
$\mathbf{C}^{KV}$ $=\mathbf{X}\bm{W}^{DKV},$
$\mathbf{K}^{R}$ $=\operatorname{RoPE}\bigl{(}\mathbf{X}\bm{W}^{KR}\bigr{)},$
E.4 Multi-matrix Factorization Attention (MFA)
Hu et al. (2024) proposed Multi-matrix Factorization Attention (MFA), which can be seen as Multi-Query Attention (MQA) with dimension of each head equals $d_{C}$, and low-rank factorized Q:
@@ E.4 Multi-matrix Factorization Attention (MFA)
Appendix F Other Variants of TPA
@@
TPA with Non-contextual B. Conversely, one may fix the token-dimension factors $\mathbf{b}^{Q}_{r},\mathbf{b}^{K}_{r},\mathbf{b}^{V}_{r}\in\mathbb{R}^{d_{h}}$ as learned parameters, while allowing $\mathbf{a}^{Q}_{r}(\mathbf{x}_{t}),\mathbf{a}^{K}_{r}(\mathbf{x}_{t}),\mathbf{a}^{V}_{r}(\mathbf{x}_{t})$ to adapt to $\mathbf{x}_{t}$. For keys:
$\mathbf{K}_{t}=\frac{1}{R_{K}}\sum_{r=1}^{R_{K}}\mathbf{a}^{K}_{r}(\mathbf{x}_{t})\otimes\mathbf{b}^{K}_{r},$
and similarly for values. This arrangement is effective if the token-dimension structure remains mostly uniform across the sequence, while the head-dimension factors capture context.
TPA KV Only. One can preserve a standard query mapping,
$\mathbf{Q}_{t}=\bm{W}^{Q}\,\mathbf{x}_{t}\in\mathbb{R}^{h\times d_{h}},$
and factorize only the keys and values. This leaves the query projection as the original linear transformation while reducing memory usage via factorized KV caching.
@@
TPA KV with Shared $\mathbf{B}$. Another variant is to share the token-dimension factors of keys and values:
$\mathbf{b}^{K}_{r}(\mathbf{x}_{t})=\mathbf{b}^{V}_{r}(\mathbf{x}_{t}),$
lowering parameter counts and the KV cache footprint. While it constrains $\mathbf{K}$ and $\mathbf{V}$ to be formed from the same token basis, it can still perform well and provide additional memory savings.
Nonlinear Head Factors. Rather than applying purely linear mappings to the head-dimension factors $\mathbf{a}^{Q}_{r},\mathbf{a}^{K}_{r},\mathbf{a}^{V}_{r}$, one may introduce element-wise nonlinearities such as $\sigma(\cdot)$ or $\operatorname{softmax}(\cdot)$. This effectively yields a Mixture of Heads Attention (MoH Attention), where each component becomes a learned mixture weight modulated by the nonlinearity.
Discussion. These variants illustrate TPA’s versatility in balancing memory cost, computational overhead, and representation power. By choosing which dimensions (heads or tokens) remain contextual and adjusting ranks $(R_{Q},R_{K},R_{V})$, TPA unifies multiple existing attention mechanisms—such as MHA, MQA, and GQA—under one framework, while potentially reducing the KV cache size by an order of magnitude during autoregressive inference.
Appendix G More on Experiments
G.1 Experimental Settings
@@
We list the main architecture hyper-parameters and training devices in Table 4. We fix $d_{h}=64$ for all the models. Moreover, we fix the number of KV heads with 2 for GQA models; $d_{h}^{R}=32$ for MLA models; and $R_{k}=R_{v}=2$, $R_{q}=6$ for TPA and TPA-KV only models. Other hyper-parameters are listed in Table 5.
@@
Model Size #Param Devices Micro BS. GAS. #Layer $d_{\text{model}}$
@@
G.2 Additional Experimental Results
G.2.1 Perplexity Curves
@@
G.2.2 Ablation Study on Different Ranks
Figure 5 shows the training loss, validation loss, and validation perplexity curves of XL-size (1.5B) T6 models with different ranks trained on the FineWeb-Edu 100B dataset, and the evaluation results are displayed in Table 7. It can be observed that an increase in rank can improve the performance of large language models.
@@ G.2.2 Ablation Study on Different Ranks
Figure 5: The training loss, validation loss and validation perplexity curves of XL-size (1.5B) T6 models with different ranks on the FineWeb-Edu 100B dataset.
G.2.3 0-shot Evaluation with lm-evaluation-harness
For the evaluation, We show the 0-shot performances with lm-evaluation-harness for small-size (124M) and XL-size (1.5B) models in Tables 6 and 7.
Table 6: The evaluation results of small models with different attention mechanisms pre-trained using FineWeb-Edu 100B dataset (0-shot with lm-evaluation-harness). The best scores in each column are bolded. Abbreviations: HellaSw. = HellaSwag, W.G. = WinoGrande.
@@
Table 7: The evaluation results of XL models with different attention mechanisms pre-trained using the FineWeb-Edu 100B dataset (0-shot with lm-evaluation-harness). The best scores in each column are bolded. Abbreviations: HellaSw. = HellaSwag, W.G. = WinoGrande. If not specified, TPA and TPA-KVonly set $R_{K}=R_{V}=2$.
@@
G.2.4 2-shot Evaluation with lm-evaluation-harness
We also show 2-shot performances in Tables 8910 and 11.
Table 8: The evaluation results of small models with different attention mechanisms on FineWeb-Edu 100B dataset (2-shot with lm-evaluation-harness). The best scores in each column are bolded. Abbreviations: HellaSw. = HellaSwag, W.G. = WinoGrande.
@@
Table 9: The evaluation results of medium models with different attention mechanisms pre-trained using FineWeb-Edu 100B dataset (2-shot with lm-evaluation-harness). The best scores in each column are bolded. Abbreviations: HellaSw. = HellaSwag, W.G. = WinoGrande.
@@
Table 10: The evaluation results of large models with different attention mechanisms pre-trained using the FineWeb-Edu 100B dataset (2-shot with lm-evaluation-harness). The best scores in each column are bolded. Abbreviations: HellaSwag = HellaSwag, WG = WinoGrande.
Method ARC-E ARC-C BoolQ HellaSwag OBQA PIQA WG MMLU SciQ Avg.
@@
Table 11: The evaluation results of XL models with different attention mechanisms pre-trained using the FineWeb-Edu 100B dataset (2-shot with lm-evaluation-harness). The best scores in each column are bolded. Abbreviations: HellaSwag = HellaSwag, WG = WinoGrande. If not specified, We set $R_{K}=R_{V}=2$ for TPA and TPA-KVonly.
Method ARC-E ARC-C BoolQ HellaSwag OBQA PIQA WG MMLU SciQ Avg.

### v4

@@ H.1 Multi-Query Attention (MQA)
Multi-Query Attention (MQA) [45] significantly reduces memory usage, particularly for the KV cache, by sharing a single key and value projection across all attention heads, while each head maintains a unique query projection. Given a sequence of input embeddings $\mathbf{X}\in\mathbb{R}^{T\times d_{\text{model}}}$, the query, shared key, and shared value tensors are computed as:
@@
Thus, each head $i$ uses a distinct query projection $\mathbf{Q}_{i}\in\mathbb{R}^{T\times d_{h}}$ but shares the common key $\mathbf{K}_{\text{shared}}\in\mathbb{R}^{T\times d_{h}}$ and value $\mathbf{V}_{\text{shared}}\in\mathbb{R}^{T\times d_{h}}$ tensors. The weight matrices are:
@@
By sharing key and value projections, MQA substantially reduces memory demands, especially for the KV cache during autoregressive inference. However, this comes at the cost of reduced model expressivity, as all heads must utilize the same key and value representations.
H.2 Grouped Query Attention (GQA)
Grouped Query Attention (GQA) [2] generalizes Multi-Head Attention (MHA) and MQA by dividing the total $h$ attention heads into $G$ groups. Within each group, heads share a common key and value projection, while each head maintains its own unique query projection. Formally, let $g(i)$ denote the group index for head $i\in\{1,\dots,h\}$, where $g(i)\in\{1,\dots,G\}$. The projections are:
@@ H.2 Grouped Query Attention (GQA)
Here, $\bm{W}^{K}_{g}$ and $\bm{W}^{V}_{g}$ are the shared weight matrices for group $g$, each in $\mathbb{R}^{d_{\text{model}}\times d_{h}}$, and $\bm{W}^{Q}_{i}\in\mathbb{R}^{d_{\text{model}}\times d_{h}}$ is the query weight matrix for head $i$. The complete output is again a concatenation of all heads:
@@
By varying $G$ from $1$ (equivalent to MQA) to $h$ (equivalent to MHA), GQA offers a trade-off between memory efficiency and model capacity.
H.3 Multi-head Latent Attention (MLA)
Multi-head Latent Attention (MLA), as used in DeepSeek-V2 [31] and DeepSeek-V3 [32], introduces low-rank compression for keys and values to reduce KV caching costs during inference.
@@ H.3 Multi-head Latent Attention (MLA)
Here, $\bm{W}^{DKV}\in\mathbb{R}^{d_{\text{model}}\times d_{c}}$ projects to a compressed dimension $d_{c}$, $\bm{W}^{UK}\in\mathbb{R}^{d_{c}\times(d_{h}h)}$ up-projects the compressed keys, $\bm{W}^{KR}\in\mathbb{R}^{d_{\text{model}}\times d_{h}^{R}}$ projects to a residual key component for RoPE, and $\bm{W}^{UV}\in\mathbb{R}^{d_{c}\times(d_{h}h)}$ up-projects the compressed values. $\mathbf{C}^{KV}\in\mathbb{R}^{T\times d_{c}}$ is the shared compressed KV latent (where $d_{c}\ll d_{h}h$). The RoPE transformation is applied to a separate key embedding $\mathbf{K}^{R}\in\mathbb{R}^{T\times d_{h}^{R}}$. Thus, only $\mathbf{C}^{KV}$ and $\mathbf{K}^{R}$ are cached, reducing KV memory usage while largely preserving performance compared to standard MHA [59].
@@
The weight matrices are $\bm{W}^{DQ}\in\mathbb{R}^{d_{\text{model}}\times d_{c}^{\prime}}$, $\bm{W}^{UQ}\in\mathbb{R}^{d_{c}^{\prime}\times(d_{h}h)}$, and $\bm{W}^{QR}\in\mathbb{R}^{d_{c}^{\prime}\times(d_{h}^{R}h)}$. Here, $\mathbf{C}^{Q}\in\mathbb{R}^{T\times d_{c}^{\prime}}$ (where $d_{c}^{\prime}\ll d_{h}h$) is the compressed query latent. The final query $\mathbf{Q}_{i}$ for each head, formed by concatenating $\mathbf{Q}_{i}^{C}$ and $\mathbf{Q}_{i}^{R}$, has a dimension of $d_{h}+d_{h}^{R}$.
@@
where $\mathbf{V}_{i}$ is typically $\mathbf{V}_{i}^{C}$ as no residual value component is explicitly defined, and $\bm{W}^{O}\in\mathbb{R}^{(d_{h}h)\times d_{\text{model}}}$ is the output projection.
During inference, $\mathbf{C}^{KV}$ and $\mathbf{K}^{R}$ are cached to accelerate decoding. In detail, if RoPE were ignored for the compressed components, the inner product $\mathbf{q}_{t,i}^{\top}\mathbf{k}_{s,i}$ (where $\mathbf{q}_{t,i},\mathbf{k}_{s,i}\in\mathbb{R}^{d_{h}}$) of the $i$-th head between $t$-th token query and $s$-th token key could be calculated using the current hidden state $\mathbf{x}_{t}\in\mathbb{R}^{d_{\text{model}}}$ and the cached latent state $\mathbf{c}_{s}^{KV}\in\mathbb{R}^{d_{c}}$ for the $s$-th token:
$\mathbf{q}_{t,i}^{\top}\mathbf{k}_{s,i}$ $=[(\bm{W}^{UQ}_{i})^{\top}(\bm{W}^{DQ}_{i})^{\top}\mathbf{x}_{t}]^{\top}[(\bm{W}^{UK}_{i})^{\top}\mathbf{c}^{KV}_{s}]$ (H.1)
$=\mathbf{x}_{t}^{\top}[\bm{W}^{DQ}_{i}\bm{W}^{UQ}_{i}(\bm{W}^{UK}_{i})^{\top}]\mathbf{c}^{KV}_{s},$ (H.2)
where $\bm{W}_{i}^{(\cdot)}$ denotes the $i$-th head’s portion of the respective weight matrix.
The term $[\bm{W}^{DQ}_{i}\bm{W}^{UQ}_{i}(\bm{W}^{UK}_{i})^{\top}]$ could be pre-computed for faster decoding. However, as noted by [50], this pre-computation strategy is not directly compatible with RoPE if RoPE were applied to these compressed representations. RoPE applies a rotation matrix $\mathbf{T}_{t}\in\mathbb{R}^{d_{h}\times d_{h}}$ based on position $t$ (see Section H.5), satisfying $\mathbf{T}_{t}\mathbf{T}_{s}^{\top}=\mathbf{T}_{t-s}$ (Equation H.4). If RoPE were applied to the up-projected $Q^{C}$ and $K^{C}$:
@@
Unlike Equation (H.2), acceleration by pre-computing the term $[\bm{W}^{DQ}_{i}\bm{W}^{UQ}_{i}\mathbf{T}_{t-s}(\bm{W}^{UK}_{i})^{\top}]$ is not possible because it depends on the relative position $(t-s)$ and thus varies for different $(t,s)$ pairs. To maintain RoPE compatibility while benefiting from compression, MLA introduces an additional, smaller key component $\mathbf{K}^{R}$ (and similarly $\mathbf{Q}^{R}$) to which RoPE is applied, while the main compressed components $\mathbf{K}^{C}$ and $\mathbf{V}^{C}$ (derived from $\mathbf{C}^{KV}$) remain RoPE-free. As we will demonstrate in Section 3.2 of the main paper, TPA offers a different approach to integrate RoPE efficiently with factorized attention through its tensor product formulation.
H.4 Multi-matrix Factorization Attention (MFA)
[20] proposed Multi-matrix Factorization Attention (MFA), which can be conceptualized as a variation of MQA where the shared key and value projections have a dimension $d_{c}$, and the query projection for each head is low-rank factorized:
@@ H.4 Multi-matrix Factorization Attention (MFA)
H.5 Rotary Position Embedding (RoPE)
Many recent LLMs use rotary position embedding (RoPE; 51) to encode positional information in the query/key vectors. Specifically, for a vector at position $t$, RoPE applies a rotation matrix $\mathbf{T}_{t}\in\mathbb{R}^{d\times d}$ (where $d$ is the dimension of the query/key vectors, typically $d_{h}$ per head). $\mathbf{T}_{t}$ is a block-diagonal matrix composed of $d/2$ rotation blocks of the form $\begin{pmatrix}\cos(t\theta_{j})&-\sin(t\theta_{j})\\ \sin(t\theta_{j})&\cos(t\theta_{j})\end{pmatrix}$ for $j\in\{1,\dots,d/2\}$. The frequencies $\{\theta_{j}\}$ are typically defined as $\theta_{j}=\text{base}^{-2j/d}$, with a common base like $10000$. If $\mathbf{q}_{t}\in\mathbb{R}^{d}$ is a query (or key) row vector for a specific head at position $t$, RoPE is applied as:
$\operatorname{RoPE}(\mathbf{q}_{t})\triangleq\mathbf{q}_{t}\mathbf{T}_{t}.$
A key property of RoPE is that the inner product between RoPE-transformed vectors depends only on their relative position. For a query $\mathbf{q}_{t}$ and key $\mathbf{k}_{s}$: $(\mathbf{q}_{t}\mathbf{T}_{t})(\mathbf{k}_{s}\mathbf{T}_{s})^{\top}=\mathbf{q}_{t}\mathbf{T}_{t}\mathbf{T}_{s}^{\top}\mathbf{k}_{s}^{\top}=\mathbf{q}_{t}\mathbf{T}_{t-s}\mathbf{k}_{s}^{\top}$. This relies on the property:
$\mathbf{T}_{t}\mathbf{T}_{s}^{\top}=\mathbf{T}_{t-s},$ (H.4)
which embeds relative positional information $(t-s)$ into the attention scores.
Appendix I More on TPA
Parameter Initialization for TPA Factors. We initialize the weight matrices for TPA factors, such as $\bm{W}_{r}^{a^{Q}}$, $\bm{W}_{r}^{a^{K}}$, $\bm{W}_{r}^{a^{V}}$, $\bm{W}_{r}^{b^{Q}}$, $\bm{W}_{r}^{b^{K}}$, and $\bm{W}_{r}^{b^{V}}$ (or their combined forms $\bm{W}^{a^{Q}}$, $\bm{W}^{b^{Q}}$, etc.), using Xavier initialization [14]. Specifically, each entry of a weight matrix is drawn from a uniform distribution $\mathcal{U}(-bound,bound)$, where $bound=\sqrt{6/(n_{\text{in}}+n_{\text{out}})}$. Here, $n_{\text{in}}$ and $n_{\text{out}}$ are the input and output dimensions of the respective weight matrix. This initialization strategy is chosen to help maintain the variance of activations and gradients as they propagate through the network layers, contributing to stable training.
@@ Appendix I More on TPA
TPA with Non-contextual B. In Section A.1, we have introduced TPA with non-contextual A, where head-dimension factors $\mathbf{a}_{r}^{Q},\mathbf{a}_{r}^{K},\mathbf{a}_{r}^{V}\in\mathbb{R}^{h}$ are fixed. Conversely, one may fix the token-dimension factors $\mathbf{b}^{Q}_{r},\mathbf{b}^{K}_{r},\mathbf{b}^{V}_{r}\in\mathbb{R}^{d_{h}}$ as learned parameters, while allowing $\mathbf{a}^{Q}_{r}(\mathbf{x}_{t}),\mathbf{a}^{K}_{r}(\mathbf{x}_{t}),\mathbf{a}^{V}_{r}(\mathbf{x}_{t})$ to adapt to the input token $\mathbf{x}_{t}$. The key tensor for token $t$, $\mathbf{K}_{t}\in\mathbb{R}^{h\times d_{h}}$, would then be constructed as:
$\mathbf{K}_{t}=\frac{1}{R_{K}}\sum_{r=1}^{R_{K}}\mathbf{a}^{K}_{r}(\mathbf{x}_{t})\otimes\mathbf{b}^{K}_{r}.$
A similar formulation applies to values. This configuration might be effective if the fundamental token-level features (captured by $\mathbf{b}_{r}$) are relatively stable, while their combination across heads (captured by $\mathbf{a}_{r}(\mathbf{x}_{t})$) needs to adapt to the context. Performance comparisons for TPA with non-contextual A factors versus non-contextual B factors on small and medium-sized models are presented in Tables 6, 7, 8, and 9.
Table 6: Evaluation results of small models with TPA using non-contextual A or B factors, pre-trained on FineWeb-Edu 100B dataset (0-shot with lm-evaluation-harness). Abbreviations: HellaSw. = HellaSwag, W.G. = WinoGrande.
Method ARC-E ARC-C BoolQ HellaSw. OBQA PIQA W.G. MMLU SciQ Avg.
TPA (non-ctx-B) 47.39 26.37 54.8 32.71 30.2 63.38 50.2 23.13 64.8 43.66
Table 7: Evaluation results of small models with TPA using non-contextual A or B factors, pre-trained on FineWeb-Edu 100B dataset (2-shot with lm-evaluation-harness). Abbreviations: HellaSw. = HellaSwag, W.G. = WinoGrande.
Method ARC-E ARC-C BoolQ HellaSw. OBQA PIQA W.G. MMLU SciQ Avg.
TPA (non-ctx-B) 50.8   26.96 57.65 32.4 29.4 63.22 49.57 23.96 66.4 44.48
Table 8: Evaluation results of medium models with TPA using non-contextual A or B factors, pre-trained on FineWeb-Edu 100B dataset (0-shot with lm-evaluation-harness). Abbreviations: HellaSw. = HellaSwag, W.G. = WinoGrande.
Method ARC-E ARC-C BoolQ HellaSw. OBQA PIQA W.G. MMLU SciQ Avg.
TPA (non-ctx-B) 55.43 29.69 58.32 40.77 34.40 66.92 51.38 25.66 71.10 48.19
Table 9: Evaluation results of medium models with TPA using non-contextual A or B factors, pre-trained on FineWeb-Edu 100B dataset (2-shot with lm-evaluation-harness). Abbreviations: HellaSw. = HellaSwag, W.G. = WinoGrande.
Method ARC-E ARC-C BoolQ HellaSw. OBQA PIQA W.G. MMLU SciQ Avg.
TPA (non-ctx-B) 61.20 30.20 55.93 40.45 34.40 68.23 51.78 26.11 78.10 49.60
TPA KV Only. A simpler variant involves using a standard linear projection for queries,
$\mathbf{Q}_{t}=\mathbf{W}^{Q}\mathbf{x}_{t}\in\mathbb{R}^{h\times d_{h}},$
and factorize only the key and value tensors ($\mathbf{K}_{t},\mathbf{V}_{t}$). This approach, termed TPA-KVonly, maintains the standard query projection mechanism but still achieves significant KV cache reduction through factorized key and value representations.
@@
TPA KV with Shared $\mathbf{B}$. Further parameter reduction can be achieved by sharing the token-dimension factors $\mathbf{b}_{r}$ between keys and values:
[…]
This sharing reduces both parameter count and the KV cache footprint. Although it constrains $\mathbf{K}_{t}$ and $\mathbf{V}_{t}$ to be constructed from the same token-level basis vectors, this variant can still offer strong performance with additional memory savings.
Nonlinear Head Factors. Instead of using purely linear transformations to derive the contextual head-dimension factors $\mathbf{a}^{Q}_{r}(\mathbf{x}_{t}),\mathbf{a}^{K}_{r}(\mathbf{x}_{t}),\mathbf{a}^{V}_{r}(\mathbf{x}_{t})$, one can introduce element-wise nonlinearities (e.g., sigmoid $\sigma(\cdot)$ or softmax). Applying softmax, for instance, to the coefficients that generate $\mathbf{a}_{r}(\mathbf{x}_{t})$ could be interpreted as a form of Mixture-of-Heads, where the network learns to dynamically weight different head configurations based on the input context.
Discussion. These variants highlight the flexibility of the TPA framework, allowing for different trade-offs between memory efficiency, computational cost, and model expressiveness. By carefully choosing which factor components (head-dimension or token-dimension) are contextual versus non-contextual, and by adjusting the ranks $(R_{Q},R_{K},R_{V})$, TPA can not only unify existing mechanisms like MHA, MQA, and GQA but also significantly reduce KV cache size—potentially by an order of magnitude—during autoregressive inference.
Appendix J More on Experiments
J.1 Experimental Settings
@@
We list the main architecture hyper-parameters and training devices in Table 10. For all models, the head dimension $d_{h}$ is fixed at 64. Specific architectural choices include: 2 KV heads for GQA models; a residual key dimension $d_{h}^{R}=32$ for MLA models; and ranks $R_{K}=R_{V}=2$ and $R_{Q}=6$ for TPA and TPA-KVonly models, unless otherwise specified. Other relevant hyper-parameters are listed in Table 11.
Training Setup Details. We follow the nanoGPT training configuration [23]. In particular, we use the AdamW [34] optimizer with $(\beta_{1},\beta_{2})=(0.9,0.95)$, a weight decay of $0.1$, and gradient clipping at $1.0$. We follow the same setting as nanoGPT that the learning rate is managed by a cosine annealing scheduler [35] with $2{,}000$ warmup steps and a (total) global batch size of $480$. For the small, medium, large and XL models, we set maximum learning rates of $6\times 10^{-4}$, $3\times 10^{-4}$, $2\times 10^{-4}$, and $1\times 10^{-4}$ (respectively), and minimum learning rates of $3\times 10^{-5}$, $6\times 10^{-5}$, $1\times 10^{-5}$, and $1\times 10^{-5}$ (respectively).
@@
Model Size Parameters Devices Micro BS. GAS. #Layers $d_{\text{model}}$
@@
J.2 Additional Experimental Results
J.2.1 Perplexity Curves
@@
J.2.2 Ablation Study on Different Ranks
Figure 10 illustrates the training loss, validation loss, and validation perplexity for XL-sized (1.5B parameters) TPA models with varying key/value ranks ($R_{K}=R_{V}=R$, as indicated in the figure legend), trained on the FineWeb-Edu 100B dataset. Corresponding 0-shot evaluation results are presented in Table 13 (rows for TPA-KVonly with different $R_{K,V}$). These results indicate that increasing the ranks for key and value factorizations generally improves the performance of the TPA models.
@@
Figure 10: The training loss, validation loss and validation perplexity curves of XL-size (1.5B) TPA models with different key/value ranks ($R_{K}=R_{V}=R$) on the FineWeb-Edu 100B dataset.
J.2.3 0-shot Evaluation with lm-evaluation-harness
We present 0-shot evaluation results using the lm-evaluation-harness for small (124M parameters) and XL (1.5B parameters) models in Tables 12 and 13, respectively.
Table 12: Evaluation results of small models (124M) with different attention mechanisms, pre-trained on FineWeb-Edu 100B dataset (0-shot with lm-evaluation-harness). The best scores in each column are bolded. Abbreviations: HellaSw. = HellaSwag, W.G. = WinoGrande.
@@
Table 13: Evaluation results of XL models (1.5B) with different attention mechanisms, pre-trained on the FineWeb-Edu 100B dataset (0-shot with lm-evaluation-harness). The best scores in each column are bolded. Abbreviations: HellaSw. = HellaSwag, W.G. = WinoGrande. If not specified, TPA and TPA-KVonly models use $R_{K}=R_{V}=2$.
@@
J.2.4 2-shot Evaluation with lm-evaluation-harness
Similarly, 2-shot evaluation results are provided in Tables 14 (Small), 15 (Medium), 16 (Large), and 17 (XL).
Table 14: Evaluation results of small models (124M) with different attention mechanisms, pre-trained on FineWeb-Edu 100B dataset (2-shot with lm-evaluation-harness). The best scores in each column are bolded. Abbreviations: HellaSw. = HellaSwag, W.G. = WinoGrande.
@@
Table 15: Evaluation results of medium models (353M) with different attention mechanisms, pre-trained on FineWeb-Edu 100B dataset (2-shot with lm-evaluation-harness, default LR $6\times 10^{-4}$). The best scores in each column are bolded. Abbreviations: HellaSw. = HellaSwag, W.G. = WinoGrande.
@@
Table 16: Evaluation results of large models (772M) with different attention mechanisms, pre-trained on the FineWeb-Edu 100B dataset (2-shot with lm-evaluation-harness). The best scores in each column are bolded. Abbreviations: HellaSw. = HellaSwag, W.G. = WinoGrande.
Method ARC-E ARC-C BoolQ HellaSw. OBQA PIQA W.G. MMLU SciQ Avg.
@@
Table 17: Evaluation results of XL models (1.5B) with different attention mechanisms, pre-trained on the FineWeb-Edu 100B dataset (2-shot with lm-evaluation-harness). The best scores in each column are bolded. Abbreviations: HellaSw. = HellaSwag, W.G. = WinoGrande. If not specified, $R_{K}=R_{V}=2$ for TPA and TPA-KVonly models.
Method ARC-E ARC-C BoolQ HellaSw. OBQA PIQA W.G. MMLU SciQ Avg.